OpenStudy (anonymous):

x4/3 − 8x2/3 + 15 = 0

OpenStudy (anonymous):

That is x^4/3 - 8x^2/3 +15 =0

OpenStudy (lgbasallote):

let x^2/3 = a a^2 - 8a + 15 = 0 (a- 5)(a - 3) = 0 a= 5 a = 3 x^2/3 = 5 x^2/3 = 3 x = cuberoot 25 x = cuberoot 9